2. A charge q sits at the origin. What is the flux of E through a plane triangular surface with corners at (x- 2,0,0), (0,y-3.0) (Hint: Use Gausss law)
Answer #1

There should be 4 triangluar surfaces are arranged to make a closed surface that can act like a gaussian suface enclosed charge q.

Total Flux = q/ϵ

or Flux through 4 surfaces = q/ϵ

Flux through each triangle = q/4ϵ

(0,0,0